OpenStudy (anonymous):

solve x + 7 + 6x -3 will give medal!

OpenStudy (paki):

combine the like terms... and solve then....

OpenStudy (anonymous):

according to the thing i am doing you can. here are the choices. A. -5 B. -15 C. 2 D. -10

OpenStudy (mrnood):

This is not an equation - so you cannot solve it to give a number. Please check your post - have you missed an = sign somewhere?

OpenStudy (blank ):

Are there any exponents? I agree with @MrNood ^

OpenStudy (anonymous):

nope, this is right. i guess i am not the only one stumped on this question. lol

OpenStudy (anonymous):

oh! i got it. lol. sorry guys. it is x + 7 = 6x -3

OpenStudy (anonymous):

lol, i'm dumb

OpenStudy (blank ):

x=2

OpenStudy (blank ):

x+7=6x-3 Move all terms not containing x to the right-hand side of the equation. -5x+7=-3 Move all terms not containing x to the right-hand side of the equation. -5x=-10 Divide each term in the equation by -5. x=2 Hope this helps :)

OpenStudy (anonymous):

lol, thanks

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (mrnood):

@Blank  It probably DOESN'T help Once he had the equation correct maybe he could solve the equation himself